# Toward a 2D Local Implementation of Quantum LDPC Codes

### v1

@@ I Introduction
The surface code, despite showing promising theoretical and experimental performance [1, 2, 3, 4, 5, 6], is poorly suited to large-scale fault-tolerant quantum computation due to its large qubit overhead [4, 7, 8]. As a result, there has been much effort on the development of high-rate quantum low-density parity-check (qLDPC) codes [9]. As these codes can encode multiple logical qubits, the required resources are reduced, in some instances, to a constant [10].
@@
One of the main drawbacks of these high-rate qLDPC codes is that many long-range connections are needed to implement their syndrome extraction circuits [11, 12, 13, 14]. This is a pressing issue for architectures such as superconducting qubits. There, many of the current designs only allow two-qubit gates to be performed between qubits that are 2D nearest neighbors, in which case implementing these long-range entangling gates incurs significant overhead [15, 16]. Several recent proposals have attempted to alleviate this overhead by taking advantage of more complex electrical wiring of the superconducting circuits [17, 18], employing code concatenation [19, 20], or using bosonic cat qubits [21]. Implementing these long-range connections is less problematic in architectures like neutral atoms, ion traps, or semiconductor spin qubits that can implement long-range gates through qubit movement [22, 23, 24, 25, 26, 27]. However, since movement adds additional complications associated with qubit decoherence, heating, and loss, it is worthwhile to consider schemes that limit the amount of movement. In the extreme case, one can consider qubits that are fixed in space and solely use local interactions to perform medium-range entangling gates. Such studies provide additional insight into the tradeoffs associated with engineering long-range connectivity through qubit motion or more complex electrical wiring.
@@
In this paper, we present an approach to qLDPC codes that works without qubit motion or long-range couplers, inspired by the so-called stacked model [13, 28]. In this model, we assume that the high-rate qLDPC codes of interest have the property that after embedding the code into $\mathbb{Z}^{2}$, the majority of the stabilizer generators are local; that is, their qubits are contained within a ball of constant radius. We claim that most of the work required to perform the syndrome extraction circuit with 2D local gates comes from measuring the few nonlocal generators, so measuring these generators less frequently has the potential to significantly reduce the time overhead, ideally at only a minor cost to the error correction performance of the code. It was shown in Ref. [28] that, for quantum expander codes [29], neglecting to measure a large percentage of generators could be reasonably tolerated; however, it is unclear whether such codes lend themselves well to physical implementations.
@@ II.2 Architecture
In this work, we consider an architecture where qubits are located on the vertices of an $M\times M$ grid, where $M=\Theta(\sqrt{n})$. As is natural for current superconducting quantum computing platforms, we assume that two-qubit gates can only be performed between neighboring qubits on the grid. Any two-qubit gate which interacts qubits that are not neighboring is considered a long-range gate. Circuits that do not have access to long-range gates are called 2D local circuits, and architectures that are restricted to these circuits are called 2D local architectures. This definition generalizes to architectures based on graphs other than the grid: given a graph $G=(V,E)$ with data qubits located on the vertices, the only allowed two-qubit gates are those between qubits $u,v\in V$ that share an edge $(u,v)\in E$. Similar restrictions arise if we disallow the slow movement of atoms in neutral-atom devices, in which case the only available two-qubit gates are those performed through Rydberg-Rydberg interactions. This leads to an architecture that can perform entangling gates on qubits that are some distance $R$ away, where $R$ depends on the capabilities of the device. We do not investigate this ability in this work, but we discuss it in Section V.
@@ III Routing bounds
Let $C$ be a circuit measuring $M^{2\beta}$ commuting Pauli operators $S_{1},\ldots,S_{r}$ whose radii are greater than $M^{\gamma}$. Then for teleportation routing,
@@
In our routing model, the maximum total length of the teleportation paths in a single time step is $\Theta(M^{2})$ since only a constant number of ancillary qubits are allowed, and there are $\Theta(M^{2})$ edges in the grid graph. The cost of measuring an operator of size $\Omega(M^{\gamma})$ is dominated by implementing the long-range CNOT gate between its two furthest qubits. Although this can be done in constant depth using a dynamic circuit (Fig. 2(a)), it requires a teleportation path of length $\Omega(M^{\gamma})$. Consequently, routing and measuring this one operator uses $\Omega(M^{\gamma})$ edges of the $\Theta(M^{2})$ available edges. Measuring all $M^{2\beta}$ operators thus requires $\Omega(M^{2\beta+\gamma})$ edges. In the best case, we utilize all available edges in each circuit layer, giving a circuit depth of $\Omega(M^{2\beta+\gamma-2})$. ∎
@@
Let $C$ be a circuit measuring $M^{2\beta}$ commuting operators whose radii follow a probability distribution $f(\gamma)$. Then for teleportation routing,
@@
Since we again have $\Theta(M^{2})$ edges in the grid available in each layer of the circuit, the total circuit depth is lower-bounded as in Eq. (3), as desired. ∎
@@ III.1 Greedy routing
Figure 3: Routing depth from greedy routing versus theoretical optimal routing depth as a function of code size. Code examples come from the family of BB qLDPC codes (see Section IV.1).
To benchmark the performance of the algorithm, we draw random examples of BB codes (see Section IV.1) and route the $X$-type generators while restricted to a single layer of ancillary qubits. For comparison, we compute the optimal routing depth according to Claim 2. Figure 3 shows the results of these simulations, providing evidence that the greedy routing algorithm nearly saturates Eq. (3). A constant multiple of the theory bound matches closely with the routing time of small code instances, although we begin to see the algorithm routing time deviating as we increase the block length. For code sizes of practical interest, this algorithm may be a viable option to optimize teleportation routing. Certain codes, such as the BB codes we discuss in the next section, have additional structure that allows us to manually find routing schedules that outperform those found by the greedy algorithm.
@@ IV.1 Bivariate bicycle codes
$[[n,k,d]]$ $\ell,m$ $A$ $B$ Embedding $p_{\text{mask}}$
$[[72,8,6]]$ 12,3 $x^{9}+y^{1}+y^{2}$ $1+x^{1}+x^{11}$ $\langle A_{2}A_{3}^{T},B_{1}B_{2}^{T}\rangle$ 25%
$[[90,8,6]]$ 9,5 $x^{8}+y^{4}+y$ $y^{5}+x^{8}+x^{7}$ $\langle A_{2}A_{3}^{T},B_{2}B_{1}^{T}\rangle$ 22.22%
$[[120,8,8]]$ 12,5 $x^{10}+y^{4}+y$ $1+x+x^{2}$ $\langle A_{2}A_{3}^{T},B_{1}B_{2}^{T}\rangle$ 25%
$[[150,8,8]]$ 15,5 $x^{5}+y^{2}+y^{3}$ $y^{2}+x^{7}+x^{6}$ $\langle A_{1}A_{2}^{T},B_{1}B_{3}^{T}\rangle$ 26.66%
$[[144,12,12]]$ 12,6 $x^{3}+y+y^{2}$ $y^{3}+x+x^{2}$ $\langle A_{2}A_{1}^{T},B_{1}B_{3}^{T}\rangle$ 33.33%
$[[196,12,8]]$ 14,7 $x^{6}+y^{5}+y^{6}$ $1+x^{4}+x^{13}$ $\langle A_{2}A_{3}^{T},B_{1}B_{2}^{T}\rangle$ 35.71%
Table 1: Examples of BB qLDPC codes found through a computer search. Code distances were computed using the QDistRnd GAP package [64]. The Embedding column reports the specific embedding into $\mathbb{Z}^{2}$ used for that code (see Appendix A). The $p_{\text{mask}}$ column denotes the percentage of generators that are “large”, i.e., induced by the long boundary and masked during a portion of the error correction rounds.
@@
These natural embeddings make it straightforward to search for codes where the check structure is geometrically small. While the checks are not entirely local due to the two nonlocal qubits in their support, appropriately choosing $\ell$ and $m$ can make the periodic boundary conditions induce generators that are comparatively much larger. This can be done by letting $\ell\gg m$ (or $m\gg\ell$). In the resulting generator distribution, the majority of the checks are geometrically small. In the context of the stacked model, the generators that are induced by the boundary conditions are those that are measured less frequently.
Table 1 lists BB codes found by computer search which, through simulations similar to those of Ref. [28], display good masked error correction performance. To our knowledge, the codes presented here are new, with the exception of the $[[144,12,12]]$ code, which was reported in Ref. [18]. When searching, another consideration was the percentage of generators induced by the long boundary. This percentage is listed in Table 1 as $p_{\text{mask}}$.
@@ IV.2 Syndrome extraction circuits
To implement a CNOT gate between a data qubit and a distant check qubit, we use the constant-depth circuit shown in Fig. 2(a). A number of ancilla qubits equal to the length of the CNOT gate are needed, and so qubits from the upper layer are utilized, as illustrated in Fig. 2(b). Multiple long-range CNOT gates may be performed in parallel as long as the paths act on disjoint sets of qubits. Given a set of CNOT gates to perform, an order that attempts to minimize the total depth of the circuits can be found using the greedy routing algorithm introduced in Section III.1. Alternatively, we can utilize the repeated check structure of the BB codes to manually come up with highly parallelized orderings; Fig. 4 shows an example of how the Bell pairs needed for the long-range CNOT gates (red highlighted paths) can be implemented in parallel. Also see Fig. 10. This repeated parity check structure is useful for implementing generalized bicycle codes with reconfigurable atom arrays [24] and bosonic cat qubits [21].
@@ IV.3 Space-time decoder
In reality, errors may occur at any operation in the syndrome extraction circuit, including qubit initialization, single- and two-qubit gates, measurements, and idle locations. To model this, we instead consider the standard circuit-based depolarizing noise model [3], where for each operation in the circuit, an error is introduced with some probability $p$. For example, an error arising from a CNOT gate is the gate followed by one of the possible 15 non-identity two-qubit Pauli products on the control and target qubits. Although it is possible to decode circuit-level noise using the same method as for phenomological noise, it has been shown to be advantageous to instead use a space-time circuit-level decoder [23]. Here, the goal is to guess the error at specific locations in the syndrome extraction circuit. Decoding is considered a success if the guessed errors have the same effect on the logical observables as the actual error.
The input to the space-time decoder is not the syndrome of the error, but rather the parities of the syndrome measurements between error correction rounds. In the absence of errors, the syndrome between rounds should be constant, i.e. have parity of zero. A parity of one indicates that an error occurred at some point in the previous error correction round. Following the notation of Stim [74, 73], we define the $i$th detector at time $t$ to be the parity of the syndrome of the current and previous rounds $D_{i}^{(t)}=\sigma_{i}^{(t)}\oplus\sigma_{i}^{(t-1)}$. However, in the stacked model, we have the possibility of neglecting to measure certain generators for some number of rounds, $t_{m}$. As such, detectors for these generators must compare the parities of the corresponding syndromes $t_{m}$ rounds apart, $D_{i}^{(t)}=\sigma_{i}^{(t)}\oplus\sigma_{i}^{(t-t_{m})}$. Each detector allows us to determine whether errors have occurred in a specific detecting region [73] of the circuit. Figure 5(a) shows a simple example of a classical repetition code circuit with its associated detectors and highlighted detecting regions.
@@ IV.4 Circuit-level simulations
For the following simulations, we use Stim [74] to construct the circuits and build the space-time bipartite graph used for decoding. As such, we consider a circuit-level noise model in which errors occur independently on different circuit operations. For a physical error rate $p$—in this work we consider $p=0.1\%$—single-qubit gates have probability $p/10$ of experiencing the single-qubit depolarizing channel; two-qubit gates have probability $p$ of experiencing the two-qubit depolarizing channel; measurement results have probability $p$ of being flipped; qubit reset operations have probability $p/10$ of preparing the |$\ket{1}$ state instead of the |$\ket{0}$ state; and idle qubits experience a depolarizing channel with probability $p/50$. This last condition on the idle qubit error rate is perhaps somewhat optimistic. We comment on this assumption in Section V.
@@
For ease of implementation, we first separately perform circuit-level simulations of the entanglement purification protocol. The simulation consists of implementing two noisy long-range Bell pairs using a circuit similar to that depicted in Fig. 2(a) and then performing Bennett et al.’s entanglement purification protocol on the two pairs. In this simplest version of the protocol, failures are not reattempted, and only a single donor Bell pair is used. Simulating the protocol many times allows us to estimate the probability that the purification protocol succeeds and, if so, the fidelity of the purified Bell pair. Fig. 6(a) displays the results of these simulations for long-range Bell pairs of different lengths under the circuit-level error model described above. During syndrome extraction, if the entanglement purification protocol for any of the long-range CNOT gates fails, we mask the corresponding generator instead of reattempting the purifications. We can then estimate the probability that the syndrome of a generator is available, that is, all of the required purifications succeed. If the purifications do succeed, then we can also estimate the error rate of the resulting long-range CNOT gate from the fidelity of the Bell pair. In the full circuit, we then implement a direct CNOT with this error rate to represent the entire procedure. Fig. 6(b) illustrates what this means practically: assuming the long-range generators are unmasked every five rounds, the first four rounds have these long-range generators masked (hatched fill). Additionally, due to failures of the entanglement purification protocol, some short-range generators are also masked, even though we had planned for them to always be available. In the fifth round, the long-range generators are unmasked and attempted to be measured, but only if purifications succeed can we actually obtain their syndromes. Note that with this simple purification scheme, the long-range generators are less likely to succeed, since the necessary Bell pairs are between more distant qubits and more prone to failure.
@@
For the full error correction protocol, we begin each circuit with a single noiseless round to initialize the logical subspace. We then perform $t$ noisy error correction rounds using the syndrome extraction circuits defined in Section IV.2. As the short-range generators are easier to measure, we attempt to measure them every round, whereas the costly, long-range generators are unmasked and attempted every five rounds. As described above, we additionally mask both the short- and long-range generators with probability equal to that of at least one of required purifications failing. In the cases where all purifications succeed, we apply the two-qubit depolarizing channel after each CNOT gate with an error rate equal to that of a long-range CNOT gate performed using a Bell pair of the appropriate distance. Idling error rates are estimated using the number of steps needed to route and purify the source and donor Bell pairs for a given set of generators (see Fig. 10). As each step consists of Bell pair generation, purification, and implementation of the long-range CNOT gate, the actual circuit depth is 10×10$\times$10 × greater. To represent idling errors, a depolarizing channel is applied at the beginning of each error correction round to every qubit with probability equal to the total circuit depth times the idle error rate. Additionally, a depolarizing channel is applied to every qubit with probability $p=0.1\%$ at the beginning of each round. Before measuring the logical observables, we noiselessly extract the full syndrome one last time. The corresponding space-time bipartite graph is then generated, and the errors are sampled and decoded. In this work, we use a decoder based on belief propagation and ordered-statistics decoding (BP-OSD) [75, 76, 77], which consists of the min-sum BP decoder followed by an order-10 combination-sweep OSD postprocessing step.

### v2

@@ I Introduction
The surface code, despite showing promising theoretical and experimental performance [1, 2, 3, 4, 5, 6], is poorly suited to large-scale fault-tolerant quantum computation due to its large qubit overhead [4, 7, 8]. As a result, there has been much effort on the development of high-rate quantum low-density parity-check (qLDPC) codes [9]. As these codes can encode multiple logical qubits, the required space resources are reduced, in some instances, to a constant [10].
@@
One of the main drawbacks of these high-rate qLDPC codes is that many long-range connections are needed to implement their syndrome extraction circuits [11, 12, 13, 14]. This is a pressing issue for architectures such as superconducting qubits. There, many of the current designs only allow two-qubit gates to be performed between qubits that are 2D nearest neighbors, in which case implementing these long-range entangling gates incurs significant overhead [15, 16]. Several recent proposals have attempted to alleviate this overhead by taking advantage of more complex electrical wiring of the superconducting circuits [17, 18], employing code concatenation [19, 20], or using bosonic cat qubits [21]. Implementing these long-range connections is less problematic in architectures like neutral atoms, ion traps, or semiconductor spin qubits that can implement long-range gates through qubit movement [22, 23, 24, 25, 26, 27, 28]. However, since movement adds additional complications associated with qubit decoherence, heating, and loss, it is worthwhile to consider schemes that limit the amount of movement. In the extreme case, one can consider qubits that are fixed in space and solely use local interactions to perform entangling gates. Such studies provide additional insight into the tradeoffs associated with engineering long-range connectivity through qubit motion or more complex electrical wiring.
@@
In this paper, we present an approach to qLDPC codes that works without qubit motion or long-range couplers, inspired by the so-called stacked model [13, 29]. In this model, we assume that the high-rate qLDPC codes of interest have the property that after embedding the code into $\mathbb{Z}^{2}$, the majority of the stabilizer generators are local; that is, their qubits are contained within a ball of constant radius. We claim that most of the work required to perform the syndrome extraction circuit with 2D local gates comes from measuring the few nonlocal generators, so measuring these generators less frequently has the potential to significantly reduce the time overhead, ideally at only a minor cost to the error correction performance of the code. It was shown in Ref. [29] that, for quantum expander codes [30], neglecting to measure a large percentage of generators could be reasonably tolerated; however, the model considered there was narrow in scope, considering only a phenomenological noise model and neglecting the problem of embedding the codes. It is therefore unclear whether such codes lend themselves well to physical implementations. Nonetheless, their results on partial error correction were optimistic and motivated the investigation of the more realistic architecture developed in this work.
@@
Our work stands as an alternative architecture that may be more practical for near-term quantum computers without the ability to move qubits. As such, it is incomparable to schemes such as Ref. [23, 24] which allow for qubit movement. Several recent works have also proposed layered architectures [17, 18]; however, their motivation is in minimizing the number of crossings in the two-qubit gate connectivity. They achieve this through the use of long-range connections, the elimination of which is the main imposed constraint of our work. Refs. [19, 20] present asymptotically well performing concatenated schemes which use only local connectivity; however, the required overheads likely make them infeasible for near- and mid-term quantum computers. In particular, Ref. [20] estimates that $\sim$600 physical qubits would be needed per each logical qubit, which is an order of magnitude more than what our architecture needs to implement the $[[144,12,12]]$ Gross code [18], with $\sim$48 physical qubits per logical qubit. Most closely comparable to our work is Ref. [15], which aimed to implement quantum expander codes with local connectivity by using a similar teleportation-based scheme. Whereas they arrived at a negative result, the innovations in code choice, partial error correction, and syndrome extraction using entanglement purification presented here allow us to obtain more favorable performance. In general, our approach may be easier to implement as it only requires a bilayer architecture, local connectivity, and relatively few qubits. It, of course, also comes with challenges, which we later discuss.
@@ II.2 Architecture
In this work, we consider an architecture where qubits are located on the vertices of an $M\times M$ grid, where $M=\Theta(\sqrt{n})$. As is natural for current superconducting quantum computing platforms, we assume that two-qubit gates can only be performed between neighboring qubits on the grid. Any two-qubit gate which interacts qubits that are not neighboring is considered a long-range gate. Circuits that do not have access to long-range gates are called 2D local circuits, and architectures that are restricted to these circuits are called 2D local architectures. This definition generalizes to architectures based on graphs other than the grid: given a connectivity graph $G=(V,E)$ with data qubits located on the vertices, the only allowed two-qubit gates are those between qubits $u,v\in V$ that share an edge $(u,v)\in E$. Similar restrictions arise if we disallow the slow movement of atoms in neutral-atom devices, in which case the only available two-qubit gates are those performed through Rydberg-Rydberg interactions. This leads to an architecture that can perform entangling gates on qubits that are some distance $R$ away, where $R$ depends on the capabilities of the device. We do not investigate this ability in this work, but we discuss it in Section V.
@@ III Routing bounds
Let $C$ be a 2D local circuit measuring $M^{2\beta}$ commuting Pauli operators whose radii are greater than $M^{\gamma}$ after embedding them in an $M\times M$ grid. Then for teleportation routing,
@@
In our routing model, the maximum total length of the teleportation paths in a single time step is $O(M^{2})$ since only a constant number of ancillary qubits per data qubit are allowed, and there are $\Theta(M^{2})$ edges in the grid graph. The cost of measuring an operator of size $\Omega(M^{\gamma})$ is dominated by implementing the long-range CNOT gate between its two furthest qubits. Although this can be done in constant depth using a dynamic circuit (Fig. 2(a)), it requires a teleportation path of length $\Omega(M^{\gamma})$. Consequently, routing and measuring this one operator uses $\Omega(M^{\gamma})$ edges of the $O(M^{2})$ available edges. Measuring all $M^{2\beta}$ operators thus requires $\Omega(M^{2\beta+\gamma})$ edges. In the best case, we utilize all available edges in each circuit layer, giving a circuit depth of $\Omega(M^{2\beta+\gamma-2})$. ∎
@@
Let $C$ be a 2D local circuit measuring $M^{2\beta}$ commuting Pauli operators whose radii follow a probability distribution $f(\gamma)$ after embedding them in an $M\times M$ grid. Then for teleportation routing,
@@
Since we again have $O(M^{2})$ edges in the grid available in each layer of the circuit, the total circuit depth is lower-bounded as in Eq. (3), as desired. ∎
@@ III.1 Greedy routing
Figure 3: Depth from greedy routing versus 2.5×2.5$\times$2.5 × the theoretical optimal depth to route the $X$-type generators using a single layer of ancillary qubits. Code examples are drawn randomly from the family of BB qLDPC codes (see Section IV.1).
To benchmark the performance of the algorithm, we draw random examples of BB codes (see Section IV.1) and route the $X$-type generators while restricted to a single layer of ancillary qubits. For comparison, we compute the optimal routing depth according to Claim 2. Figure 3 shows the results of these simulations, providing evidence that the greedy routing algorithm nearly saturates Eq. (3). To obtain the constant multiple in Fig. 3, we consider the smallest eight code instances and perform a fit between the asymptotic lower bound and the depth returned from the greedy routing algorithm. This constant times the theory lower bound matches closely with the routing time of small code instances; however, we begin to see the algorithm routing depth deviating as we increase the block length, indicating non-optimal performance. For code sizes of practical interest, this algorithm may be a viable option to optimize teleportation routing. Certain codes, such as the BB codes we discuss in the next section, have additional structure that allows us to manually find routing schedules that outperform those found by the greedy algorithm.
@@ IV.1 Bivariate bicycle codes
$[[n,k,d]]$ $\ell,m$ $A$ $B$ Embedding Mask percent Routing steps
$[[72,8,6]]$ 12,3 $x^{9}+y^{1}+y^{2}$ $1+x^{1}+x^{11}$ $\langle A_{2}A_{3}^{T},B_{1}B_{2}^{T}\rangle$ 25% 11,6
$[[90,8,6]]$ 9,5 $x^{8}+y^{4}+y$ $y^{5}+x^{8}+x^{7}$ $\langle A_{2}A_{3}^{T},B_{2}B_{1}^{T}\rangle$ 22.22% 9,5
$[[120,8,8]]$ 12,5 $x^{10}+y^{4}+y$ $1+x+x^{2}$ $\langle A_{2}A_{3}^{T},B_{1}B_{2}^{T}\rangle$ 25% 11,6
$[[150,8,8]]$ 15,5 $x^{5}+y^{2}+y^{3}$ $y^{2}+x^{7}+x^{6}$ $\langle A_{1}A_{2}^{T},B_{1}B_{3}^{T}\rangle$ 26.66% 11,6
$[[144,12,12]]$ 12,6 $x^{3}+y+y^{2}$ $y^{3}+x+x^{2}$ $\langle A_{2}A_{1}^{T},B_{1}B_{3}^{T}\rangle$ 33.33% 12,8
$[[196,12,8]]$ 14,7 $x^{6}+y^{5}+y^{6}$ $1+x^{4}+x^{13}$ $\langle A_{2}A_{3}^{T},B_{1}B_{2}^{T}\rangle$ 35.71% 16,15
Table 1: Examples of BB qLDPC codes found through a computer search. Code distances were computed using the QDistRnd GAP package [65], with 1000 information sets and mindist = 0 to obtain the actual distance. The ‘Embedding’ column reports the specific embedding into $\mathbb{Z}^{2}$ used for that code (see Appendix A). The ‘Mask percent’ column denotes the percentage of generators that are “large”, i.e., induced by the long boundary and masked during a portion of the error correction rounds. The ‘Routing’ steps column indicates the number of routing rounds required to route, purify, and measure the short-range and long-range generators, respectively. Algorithm 1 was not used to determine the circuits; instead, the repeated generator structure of the BB codes allowed us to find circuits by hand. The actual circuit depth is $\times1$ greater due to the Bell pair generation (depth 6), purification (depth 2), and implementation of the long-range CNOT gate (depth 3).
@@
These natural embeddings make it straightforward to search for codes where the check structure is geometrically small. While the checks are not entirely local due to the two nonlocal qubits in their support, appropriately choosing $\ell$ and $m$ can make the periodic boundary conditions induce generators that are comparatively much larger. This can be done by letting $\ell\gg m$, as illustrated in Fig. 10. In the resulting generator distribution, the majority of the checks are geometrically small. In the context of the stacked model, the generators that are induced by the boundary conditions are those that are measured less frequently.
Table 1 lists BB codes found by computer search which, through simulations similar to those of Ref. [29], display good numerical performance. For each code, every valid embedding was simulated in a simplified version of the protocol in Section IV.4 in order to find the embedding that yielded the best masked error correction performance. Choosing an embedding determined the percentage of generators induced by the long boundary. This percentage is listed in Table 1 in the Mask percent column. To our knowledge, the codes presented here are new, with the exception of the $[[144,12,12]]$ code, which was reported in Ref. [18].
@@ IV.2 Syndrome extraction circuits
To implement a CNOT gate between a data qubit and a distant check qubit, we use the constant-depth circuit shown in Fig. 2(a). A number of ancilla qubits equal to the length of the CNOT gate are needed, and so qubits from the upper layer are utilized, as illustrated in Fig. 2(b). Multiple long-range CNOT gates may be performed in parallel as long as the paths act on disjoint sets of qubits. Given a set of CNOT gates to perform, an order that attempts to minimize the total depth of the circuits can be found using the greedy routing algorithm introduced in Section III.1. Alternatively, we can utilize the repeated check structure of the BB codes to manually come up with highly parallelized orderings; Fig. 4 shows an example of how the Bell pairs needed for the long-range CNOT gates (red highlighted paths) can be implemented in parallel. Also see Fig. 9. The ‘Routing steps’ column in Table 1 indicates the number of routing rounds required to route, purify, and measure the short-range and long-range generators, respectively, for these hand-designed orderings. This repeated parity check structure is also useful for implementing generalized bicycle codes with reconfigurable atom arrays [24] and bosonic cat qubits [21].
@@
In reality, errors may occur at any operation in the syndrome extraction circuit, including qubit initialization, single- and two-qubit gates, measurements, and idle locations. To model this, we instead consider the standard circuit-based depolarizing noise model [3], where for each operation in the circuit, an error is introduced with some probability $p$. For example, an error arising from a CNOT gate is the gate followed by one of the possible 15 non-identity two-qubit Pauli products on the control and target qubits. Although it is possible to decode circuit-level noise using the same method as for phenomological noise, it has been shown to be advantageous to instead use a space-time circuit-level decoder [23, 74]. Here, the goal is to guess the error at specific locations in the syndrome extraction circuit. Decoding is considered a success if the guessed errors have the same effect on the logical observables as the actual error.
The input to the space-time decoder is not the syndrome of the error, but rather the parities of the syndrome measurements between error correction rounds. In the absence of errors, the syndrome between rounds should be constant, i.e. have parity of zero. A parity of one indicates that an error occurred at some point in the previous error correction round. Following the notation of Stim [75, 76], we define the $i$th detector at time $t$ to be the parity of the syndrome of the current and previous rounds $D_{i}^{(t)}=\sigma_{i}^{(t)}\oplus\sigma_{i}^{(t-1)}$, where $\sigma_{i}^{(t)}$ is the $i$th bit of the syndrome at time $t$. However, in the stacked model, we have the possibility of neglecting to measure certain generators for some number of rounds, $t_{m}$. As such, detectors for these generators must compare the parities of the corresponding syndromes $t_{m}$ rounds apart, $D_{i}^{(t)}=\sigma_{i}^{(t)}\oplus\sigma_{i}^{(t-t_{m})}$. Each detector allows us to determine whether errors have occurred in a specific detecting region [76] of the circuit. Figure 5(a) shows a simple example of a classical repetition code circuit with its associated detectors and highlighted detecting regions.
@@ IV.4 Circuit-level simulations
For the following simulations, we use Stim [75] to construct the circuits and build the space-time bipartite graph used for decoding. As such, we consider a circuit-level noise model in which errors occur independently on different circuit operations. For a physical error rate $p$—in this work we consider $p=0.1\%$—single-qubit gates have probability $p/10$ of experiencing the single-qubit depolarizing channel; two-qubit gates have probability $p$ of experiencing the two-qubit depolarizing channel; measurement results have probability $p$ of being flipped; qubit reset operations have probability $p/10$ of preparing the |$\ket{1}$ state instead of the |$\ket{0}$ state; and idle qubits experience a depolarizing channel with probability $p/50$. The assumed single-qubit, two-qubit, and measurement error rates are comparable to the performance of current ion-trap [26, 77, 78] and superconducting [79] quantum computers. However, this last condition on the idle qubit error rate is somewhat optimistic and is around an order of magnitude better than the idle error seen on production devices. We comment on this assumption in Section V.
@@
For ease of implementation, we first separately perform circuit-level simulations of the entanglement purification protocol. The simulation consists of implementing two noisy long-range Bell pairs using a circuit similar to that depicted in Fig. 2(a) and then performing Bennett et al.’s entanglement purification protocol on the two pairs. In this simplest version of the protocol, failures are not reattempted, and only a single donor Bell pair is used. Simulating the protocol many times allows us to estimate the probability that the purification protocol succeeds and, if so, the fidelity of the purified Bell pair. Fig. 6(a) displays the results of these simulations for long-range Bell pairs of different lengths under the circuit-level error model described above.
During syndrome extraction, if the entanglement purification protocol for any of the long-range CNOT gates fails, we mask the corresponding generator instead of reattempting the purifications. We can then estimate the probability that the syndrome of a generator is available, that is, all the required purifications for that generator succeed. If the purifications do succeed, then we can also estimate the error rate of the resulting long-range CNOT gate from the fidelity of the Bell pair. In the full circuit, we then implement a direct CNOT with this error rate to represent the entire procedure. Fig. 6(b) illustrates what this means practically: assuming the long-range generators are unmasked every five rounds, the first four rounds have these long-range generators masked (hatched fill). Additionally, due to failures of the entanglement purification protocol, some short-range generators are also masked, even though we had planned for them to always be available. We note that these random failures are not expected to greatly impact the performance of the code, as it is unlikely that one generator will fail several rounds in a row. Thus, even if there are missed errors, they will likely be corrected when the generator does succeed in routing. In the fifth round, the long-range generators are unmasked and attempted to be measured, but only if purifications succeed can we actually obtain their syndromes. Note that with this simple purification scheme, the long-range generators are less likely to succeed, since the necessary Bell pairs are between more distant qubits and more prone to failure.
@@
For the full error correction protocol, we begin each circuit with a single noiseless round to initialize the logical subspace. We then perform $t$ noisy error correction rounds using the syndrome extraction circuits defined in Section IV.2. As the short-range generators are easier to measure, we attempt to measure them every round, whereas the costly, long-range generators are unmasked and attempted every five rounds. As described above, we additionally mask both the short- and long-range generators with probability equal to that of at least one of required purifications failing. In the cases where all purifications for a single generator succeed, we apply the two-qubit depolarizing channel after each CNOT gate with an error rate equal to that of a long-range CNOT gate performed using a Bell pair of the appropriate distance. Idling error rates are estimated using the number of steps needed to route and purify the source and donor Bell pairs for a given set of generators (see Fig. 9 and Table 1). As each step consists of Bell pair generation (depth 6), purification (depth 2), and implementation of the long-range CNOT gate (depth 3), the actual circuit depth is $\times1$ greater. To represent idling errors, a depolarizing channel is applied at the beginning of each error correction round to every qubit with probability equal to the total circuit depth times the idle error rate. Additionally, a depolarizing channel is applied to every qubit with probability $p=0.1\%$ at the beginning of each round. Before measuring the logical observables, we noiselessly extract the full syndrome one last time. The corresponding space-time bipartite graph is then generated, and the errors are sampled and decoded.
In this work, we use a decoder based on belief propagation and ordered-statistics decoding (BP-OSD) [80, 81, 82], which consists of the min-sum BP decoder followed by an order-10 combination-sweep OSD postprocessing step. Performing real time decoding using BP and higher-order OSD postprocessing may be infeasible within the fast cycle time of superconducting quantum computers; however, it was shown that good decoding performance for BB codes can be achieved while using less computationally expensive OSD parameters [83].
$[[n,k,d]]$ Qubits $\epsilon_{L}$
$[[128,8,4]]$ 248 $1.4\times 10^{-3}\pm 1.2\times 10^{-6}$
$\mathbf{[[72,8,6]]}$ 288 $1.6\times 10^{-3}\pm 3.0\times 10^{-5}$
$\mathbf{[[90,8,6]]}$ 360 $8.9\times 10^{-4}\pm 2.0\times 10^{-5}$
$[[200,8,5]]$ 392 $2.0\times 10^{-4}\pm 6.5\times 10^{-7}$
$\mathbf{[[120,8,8]]}$ 480 $1.2\times 10^{-4}\pm 2.0\times 10^{-6}$
$[[288,8,6]]$ 568 $9.5\times 10^{-5}\pm 2.5\times 10^{-7}$
$\mathbf{[[150,8,8]]}$ 600 $5.3\times 10^{-5}\pm 1.3\times 10^{-6}$
$[[392,8,7]]$ 776 $2.0\times 10^{-5}\pm 1.5\times 10^{-7}$
$\mathbf{[[144,12,12]]}$ 576 $1.6\times 10^{-4}\pm 4.6\times 10^{-6}$
$[[300,12,5]]$ 588 $3.0\times 10^{-4}\pm 9.8\times 10^{-7}$
$\mathbf{[[196,12,8]]}$ 784 $7.9\times 10^{-5}\pm 2.3\times 10^{-6}$
$[[432,12,6]]$ 852 $1.4\times 10^{-4}\pm 3.7\times 10^{-7}$
[588,12,7]][588,12,7]][ 588 , 12 , 7 ] ] 1164 $2.9\times 10^{-5}\pm 2.3\times 10^{-7}$
Table 2: Code parameters, total number of qubits used, and $\epsilon_{L}$ as extracted from Eq. (9) for the simulations described in Section IV.4. Code parameters shown in bold correspond to BB code instances. Code parameters not in bold correspond to copies of the rotated surface code.
Figure 7: Extracted logical error rate per round, $\epsilon_{L}$ as a function of the total number of qubits used (data qubits plus all ancilla qubits) for several BB and surface code instances. The data is tabulated in Table 2. We also include simulation results in the no idle error regime, as indicated by the black lines; these results are tabulated in Table 3.